OpenStudy (anonymous):

What is the average rate of change of temperature with respect to time over the evening hours from 6 PM to 10 PM? Interpret this value in a complete sentence.

OpenStudy (anonymous):

Time of Day 10 AM 12 Noon 2 PM 4 PM 6 PM 8 PM 10 PM Temperature (4072-03-01-01-00_files/i0290000.jpg) 30 35 36 36 34 30 28

OpenStudy (anonymous):

a. 4072-03-01-01-00_files/i0290001.jpg degrees per hour; The temperature drops at an average rate of 6 degrees per hour from 6 PM to 10 PM. b. 4072-03-01-01-00_files/i0290002.jpg degrees per hour; The temperature drops at an average rate of 6 degrees per hour from 6 PM to 10 PM. c. 4072-03-01-01-00_files/i0290003.jpg degrees per hour; The temperature drops at an average rate of 1.5 degrees per hour from 6 PM to 10 PM. d. 4072-03-01-01-00_files/i0290004.jpg degrees per hour; The temperature drops at an average rate of 1.5 degrees per hour from 6 PM to 10 PM.

Vocaloid (vocaloid):

The images won't load on my computer but... the average rate of change in a function f(x) from A to B can be represented by [f(B)-f(A)]/(B-A) in this case, we would use 6 PM and 10 PM as A and B, and calculate the average rate of change by using (temperature at 10PM - temperature at 6 PM)/(10-6)

Vocaloid (vocaloid):

if I'm reading the table correctly, it looks like the temperature at 10PM is 28 and the temperature at 6 PM is 34?

OpenStudy (anonymous):

yes sorry about the images i didnt know how to add a table to here

OpenStudy (anonymous):

thank you soooooooooo much lol your real mvp the answer is -6

OpenStudy (anonymous):

dang nvm it was -1.5 lmao i must of put it in the calculator wrong

Vocaloid (vocaloid):

(28-34)/(10-6) = -6/4 = -1.5 I think you forgot the denominator

OpenStudy (anonymous):

ohhhhh okay thankyou :)